OpenStudy (darkigloo):

if y=cosx^2 - sinx^2, find y'

OpenStudy (anonymous):

-2sinx-2cosx

OpenStudy (anonymous):

\[y'=2\cos x \left( -\sin x \right)-2\sin x \cos x=-4\sin x \cos x=-2\sin 2x\]

OpenStudy (ranga):

Which is squared? The x or the cosine?

OpenStudy (darkigloo):

x

OpenStudy (darkigloo):

i mean cos

OpenStudy (anonymous):

or y=cos^2x-sin^2x=cos 2x \[y=\cos 2x,y'=-2\sin 2x\]

OpenStudy (darkigloo):

Thank you! :)

OpenStudy (anonymous):

yw